1. How many ml of water are needed to dilute 35ml 4M KCl to 3M?
2. What is the molarity of a solution containing 450 ml of solution containing 45 g of
MgCO3?
3. What is the molality of a solution made by adding 6.5g of NaCl to 600g of water?
4. What are the concentrations of aluminum and sulfate in a 5.0 M solution of
aluminum sulfate?
1
Expert's answer
2021-09-15T02:40:35-0400

1)

Moles of KCl"= 4×\frac{35}{1000}=0.14moles"


Volume"=\frac{0.14}{3}=0.047l = 47ml"


2)


Moles of "MgCO_3=\frac{45}{84.3139}=0.534moles"


Molarity "=\frac{0.534×1000}{450}=1.187M"


3)

Moles of NaCl"=\frac{6.5}{58.44}=0.111moles"


Molarity = "\frac{0.111×1000}{600}=0.185M"


4)